OpenStudy (anonymous):

A laptop computer is purchased for $2100. After each year, the resale value decreases by 30% . What will the resale value be after 3 years?

OpenStudy (anonymous):

.30 * 2100=630 so 2100 - 630 = 1470. After 1 year it's worth 1470$. Year2: 1470*.30 =441 1470-441 = 1029. Thats the cost at the end of year two. Now year 3. 1029*.30 = 308.7 so 1029-308.7=720.3$ So after 3 years the computer is worth $720.3

OpenStudy (anonymous):

are you super positive this is right?

OpenStudy (anonymous):

All my work is right there.^

OpenStudy (anonymous):

ok. thank you Nathan!

OpenStudy (amistre64):

the question is rather vague; is it a straight-line depreciation? or a recurrsion process like Nathan did?

OpenStudy (anonymous):

recurrsion